# Mixed FEM for poroelasticity with stress-dependent permeabilityKhan, Lamichhane, Ruiz-Baier & Villa-Fuentes

### v1

@@ 2.1 Derivation of weak forms
$\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\,:=\,\big{\{}\boldsymbol{\tau}\in\mathbb{H}(\mathbf{div};\Omega):\,\boldsymbol{\tau}\mbox{\boldmath{$n$}}=\boldsymbol{0}{\quad\hbox{on}\quad}\Gamma_{\mathrm{N}}\big{\}},\quad\displaystyle\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\,:=\,\big{\{}{\boldsymbol{\psi}}\in\mathbf{H}(\mathop{\rm div}\nolimits;\Omega):\,{\boldsymbol{\psi}}\cdot\mbox{\boldmath{$n$}}=0{\quad\hbox{on}\quad}\Gamma_{\mathrm{N}}\big{\}}.$
We test equation (1.1) against $\boldsymbol{v}\in\mathbf{L}^{2}(\Omega)$, equation (1.9) against $\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$, impose the symmetry of $\boldsymbol{\sigma}$ weakly, test equation (1.11) against $q\in\mathrm{L}^{2}(\Omega)$, equation (1.10) against ${\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$, integrate by parts and using the boundary conditions (1.12) naturally, and then reorder the resulting equations. Then we arrive at
$\displaystyle\int_{\Omega}\mathcal{C}^{-1}\boldsymbol{\sigma}:\boldsymbol{\tau}+\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}p\,\mathop{\rm tr}\boldsymbol{\tau}+\int_{\Omega}\boldsymbol{u}\cdot\mathbf{div}\boldsymbol{\tau}+\int_{\Omega}\boldsymbol{\rho}:\boldsymbol{\tau}$ $=\displaystyle\langle\boldsymbol{\tau}\mbox{\boldmath{$n$}},\boldsymbol{u}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}$ $\quad\forall\,\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),$
$\displaystyle\int_{\Omega}\boldsymbol{v}\cdot\mathbf{div}\,\boldsymbol{\sigma}$ $=\displaystyle-\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{v}$ $\quad\forall\,\boldsymbol{v}\in\mathbf{L}^{2}(\Omega),$
$\displaystyle\int_{\Omega}\boldsymbol{\sigma}:{\boldsymbol{\eta}}$ $=0$ $\quad\forall\,{\boldsymbol{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),$
$\displaystyle\int_{\Omega}\kappa(\boldsymbol{\sigma},p)^{-1}{\boldsymbol{\varphi}}\cdot{\boldsymbol{\psi}}+\int_{\Omega}p\mathop{\rm div}\nolimits{\boldsymbol{\psi}}$ $=\displaystyle\langle{\boldsymbol{\psi}}\cdot\mbox{\boldmath{$n$}},p_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}$ $\quad\forall\,{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),$
$\displaystyle\big{(}c_{0}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}\big{)}\int_{\Omega}p\,q+\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}q\,\mathop{\rm tr}\boldsymbol{\sigma}-\int_{\Omega}q\mathop{\rm div}\nolimits{\boldsymbol{\varphi}}$ $=\displaystyle\int_{\Omega}g\,q$ $\quad\forall\,q\in\mathrm{L}^{2}(\Omega),$
@@
Introducing bilinear forms $a:\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\to\mathbb{R}$, $b:\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\to\mathbb{R}$, $c:\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{L}^{2}(\Omega)\to\mathbb{R}$, the nonlinear form $\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}:\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\to\mathbb{R}$, and the bilinear weak forms $\widetilde{b}:\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)\to\mathbb{R}$ and $\widetilde{c}:\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\Omega)\to\mathbb{R}$, defined by
@@
(2.1) $\begin{array}[]{c}\displaystyle a(\boldsymbol{\sigma},\boldsymbol{\tau}):=\int_{\Omega}\mathcal{C}^{-1}\boldsymbol{\sigma}:\boldsymbol{\tau},\quad b(\boldsymbol{\tau},(\boldsymbol{v},{\boldsymbol{\eta}})):=\int_{\Omega}\boldsymbol{v}\cdot\mathbf{div}\,\boldsymbol{\tau}+\int_{\Omega}\boldsymbol{\tau}:{\boldsymbol{\eta}},\quad c(\boldsymbol{\tau},q):=\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}q\,\mathop{\rm tr}\boldsymbol{\tau},\\[12.91663pt] \displaystyle\widetilde{a}_{\widehat{\boldsymbol{\sigma}},\widehat{p}}({\boldsymbol{\varphi}},{\boldsymbol{\psi}}):=\int_{\Omega}\kappa(\widehat{\boldsymbol{\sigma}},\widehat{p})^{-1}{\boldsymbol{\varphi}}\cdot{\boldsymbol{\psi}},\quad\widetilde{b}({\boldsymbol{\psi}},q):=\int_{\Omega}q\mathop{\rm div}\nolimits{\boldsymbol{\psi}},\quad\widetilde{c}(p,q):=\big{(}c_{0}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}\big{)}\int_{\Omega}p\,q,\end{array}$
@@
respectively, and linear functionals $H\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)^{\prime}$, $F\in(\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega))^{\prime}$, $\widetilde{H}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)^{\prime}$, $G\in\mathrm{L}^{2}(\Omega)^{\prime}$
@@
$H(\boldsymbol{\tau}):=\langle\boldsymbol{\tau}\mbox{\boldmath{$n$}},\boldsymbol{u}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}},\quad F(\boldsymbol{v},{\boldsymbol{\eta}}):=\displaystyle-\int_{\Omega}\boldsymbol{f}\cdot\boldsymbol{v},\quad\widetilde{H}({\boldsymbol{\psi}}):=\langle{\boldsymbol{\psi}}\cdot\mbox{\boldmath{$n$}},p_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}},\quad G(q):=\displaystyle-\int_{\Omega}g\,q,$
@@
we arrive at: find $(\boldsymbol{\sigma},\boldsymbol{u},\boldsymbol{\rho},{\boldsymbol{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$, such that:
@@
(2.2a) $\displaystyle a(\boldsymbol{\sigma},\boldsymbol{\tau})$ $+\quad\displaystyle b(\boldsymbol{\tau},(\boldsymbol{u},\boldsymbol{\rho}))$ $+\quad c(\boldsymbol{\tau},p)$ $=\displaystyle H(\boldsymbol{\tau})$ $\forall\,\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),$
@@
(2.2b) $\displaystyle b(\boldsymbol{\sigma},(\boldsymbol{v},{\boldsymbol{\eta}}))$ $=\displaystyle F(\boldsymbol{v},{\boldsymbol{\eta}})$ $\forall\,\boldsymbol{v}\in\mathbf{L}^{2}(\Omega),\,\forall\,{\boldsymbol{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),$
@@
(2.2c) $\displaystyle\widetilde{a}_{\boldsymbol{\sigma},p}({\boldsymbol{\varphi}},{\boldsymbol{\psi}})$ $+\quad\displaystyle\widetilde{b}({\boldsymbol{\psi}},p)$ $=\displaystyle\widetilde{H}({\boldsymbol{\psi}})$ $\forall\,{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),$
@@
(2.2d) $\displaystyle\widetilde{b}({\boldsymbol{\varphi}},q)$ $-\quad\widetilde{c}(p,q)$ $-\quad c(\boldsymbol{\sigma},q)$ $=\displaystyle G(q)$ $\forall\,q\in\mathrm{L}^{2}(\Omega).$
@@ 2.2 Stability properties and suitable inf-sup conditions
For the sake of the analysis, we allow the permeability $\kappa(\boldsymbol{\sigma},p)$ to be anisotropic but still require $\kappa(\boldsymbol{\sigma},p)^{-1}$ to be uniformly positive definite in $\mathbb{L}^{\infty}(\Omega)$ and Lipschitz continuous with respect to $p\in\mathrm{L}^{2}(\Omega)$. That is, there exist positive constants $\kappa_{1},\kappa_{2}$ such that
(2.3) $\kappa_{1}|\boldsymbol{v}|^{2}\leq\boldsymbol{v}^{\tt t}\kappa(\cdot,\cdot)^{-1}\boldsymbol{v},\qquad\|\kappa(\cdot,p_{1})^{-1}-\kappa(\cdot,p_{2})^{-1}\|_{\mathbb{L}^{\infty}(\Omega)}\leq\kappa_{2}\|p_{1}-p_{2}\|_{0,\Omega},$
for all $\boldsymbol{v}\in\mathbb{R}^{d}\setminus\{\boldsymbol{0}\}$, and for all $p_{1},p_{2}\in\mathrm{L}^{2}(\Omega)$.
@@
(2.4a) $\big{|}a(\boldsymbol{\sigma},\boldsymbol{\tau})\big{|}\leq\dfrac{1}{\mu}\|\boldsymbol{\sigma}\|_{\mathbf{div};\Omega}\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega},\qquad\big{|}b(\boldsymbol{\tau},(\boldsymbol{v},{\boldsymbol{\eta}}))\big{|}\leq\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}(\|\boldsymbol{v}\|_{0,\Omega}+\|{\boldsymbol{\eta}}\|_{0,\Omega}),$
(2.4b) $\big{|}c(\boldsymbol{\tau},q)\big{|}\leq\gamma\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}\|q\|_{0,\Omega},$
(2.4c) $\big{|}\widetilde{b}({\boldsymbol{\psi}},q)\big{|}\leq\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}\|q\|_{0,\Omega},\qquad\big{|}\widetilde{c}(p,q)\big{|}\leq\widetilde{\gamma}\|p\|_{0,\Omega}\|q\|_{0,\Omega},$
@@
(2.5) $\gamma:=\dfrac{\alpha\sqrt{d}}{d\lambda+2\mu}{\quad\hbox{and}\quad}\widetilde{\gamma}:=c_{0}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}.$
@@
$\big{|}H(\boldsymbol{\tau})\big{|}\leq\|\boldsymbol{u}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega},\qquad\big{|}F(\boldsymbol{v},{\boldsymbol{\eta}})\big{|}\leq\|\boldsymbol{f}\|_{0,\Omega}\|\boldsymbol{v}\|_{0,\Omega}\leq\|\boldsymbol{f}\|_{0,\Omega}(\|\boldsymbol{v}\|_{0,\Omega}+\|{\boldsymbol{\eta}}\|_{0,\Omega}),$
$\big{|}\widetilde{H}({\boldsymbol{\psi}})\big{|}\leq\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}\|{\boldsymbol{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega},\qquad\big{|}G(q)\big{|}\leq\|g\|_{0,\Omega}\|q\|_{0,\Omega}.$
@@
(2.6a) $\mathbb{V}$ =$=\displaystyle\Big{\{}\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega):\quad\mathbf{div}\,\boldsymbol{\tau}=\boldsymbol{0}{\quad\hbox{and}\quad}\boldsymbol{\tau}=\boldsymbol{\tau}^{\tt t}{\quad\hbox{in}\quad}\Omega\Big{\}},$
(2.6b) $\mathbf{V}$ =$=\Big{\{}{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega):\quad\mathop{\rm div}\nolimits\,{\boldsymbol{\psi}}=0\quad\text{in }\,\Omega\Big{\}}.$
@@
(2.7) $a(\boldsymbol{\tau},\boldsymbol{\tau})\geq c_{a}\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}^{2}\quad\forall\,\boldsymbol{\tau}\in\mathbb{V}.$
@@
(2.8a) $\sup_{\boldsymbol{0}\neq\boldsymbol{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)}\frac{b(\boldsymbol{\tau},(\boldsymbol{v},{\boldsymbol{\eta}}))}{\|\boldsymbol{\tau}\|_{\mathbf{div};\Omega}}$ $\geq\beta(\|\boldsymbol{v}\|_{0,\Omega}+\|{\boldsymbol{\eta}}\|_{0,\Omega})\quad\forall\,(\boldsymbol{v},{\boldsymbol{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),$
(2.8b) $\displaystyle\sup_{\boldsymbol{0}\neq{\boldsymbol{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)}\frac{\widetilde{b}({\boldsymbol{\psi}},q)}{\|\psi\|_{\mathop{\rm div}\nolimits;\Omega}}$ $\geq\widetilde{\beta}\|q\|_{0,\Omega}\quad\forall\,q\in\mathrm{L}^{2}(\Omega).$
@@
(2.9) $\widetilde{c}(q,q)\geq\widetilde{\gamma}\,\|q\|_{0,\Omega}^{2}.$

### v2

@@ 2.1 Derivation of weak forms
$\displaystyle\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$ :=$:=\,\big{\{}\bm{\tau}\in\mathbb{H}(\mathbf{div};\Omega):\,\bm{\tau}\mbox{\boldmath{$n$}}=\bm{0}{\quad\hbox{on}\quad}\Gamma_{\mathrm{N}}\big{\}},$
$\displaystyle\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$ :=$:=\,\big{\{}{\bm{\psi}}\in\mathbf{H}(\mathop{\rm div}\nolimits;\Omega):\,{\bm{\psi}}\cdot\mbox{\boldmath{$n$}}=0{\quad\hbox{on}\quad}\Gamma_{\mathrm{N}}\big{\}}.$
We test equation (1.1) against $\mathbf{v}\in\mathbf{L}^{2}(\Omega)$, equation (1.9) against $\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)$, impose the symmetry of $\bm{\sigma}$ weakly, test equation (1.11) against $q\in\mathrm{L}^{2}(\Omega)$, equation (1.10) against ${\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)$, integrate by parts and use the boundary conditions (1.12) naturally, and then reorder the resulting equations. Then we arrive at
$\displaystyle\int_{\Omega}\mathcal{C}^{-1}\bm{\sigma}:\bm{\tau}+\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}p\,\mathop{\rm tr}\bm{\tau}+\int_{\Omega}{\bm{u}}\cdot\mathbf{div}\bm{\tau}+\int_{\Omega}\bm{\rho}:\bm{\tau}$ $=\displaystyle\langle\bm{\tau}\mbox{\boldmath{$n$}},{\bm{u}}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}$ $\quad\forall\,\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),$
$\displaystyle\int_{\Omega}\mathbf{v}\cdot\mathbf{div}\,\bm{\sigma}$ $=\displaystyle-\int_{\Omega}\bm{f}\cdot\mathbf{v}$ $\quad\forall\,\mathbf{v}\in\mathbf{L}^{2}(\Omega),$
$\displaystyle\int_{\Omega}\bm{\sigma}:{\bm{\eta}}$ $=0$ $\quad\forall\,{\bm{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),$
$\displaystyle\int_{\Omega}{\bm{\kappa}}(\bm{\sigma},p)^{-1}{\bm{\varphi}}\cdot{\bm{\psi}}+\int_{\Omega}p\mathop{\rm div}\nolimits{\bm{\psi}}$ $=\displaystyle\langle{\bm{\psi}}\cdot\mbox{\boldmath{$n$}},p_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}}$ $\quad\forall\,{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),$
$\displaystyle\big{(}c_{0}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}\big{)}\int_{\Omega}p\,q+\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}q\,\mathop{\rm tr}\bm{\sigma}-\int_{\Omega}q\mathop{\rm div}\nolimits{\bm{\varphi}}$ $=\displaystyle\int_{\Omega}g\,q$ $\quad\forall\,q\in\mathrm{L}^{2}(\Omega),$
@@
Next we proceed to introduce the bilinear forms $a:\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\to\mathbb{R}$, $b:{\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times[\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)]\to\mathbb{R}}$, $c:\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathrm{L}^{2}(\Omega)\to\mathbb{R}$, the nonlinear form $\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}:\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\to\mathbb{R}$, and the bilinear forms $\widetilde{b}:\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)\to\mathbb{R}$ and $\widetilde{c}:\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\Omega)\to\mathbb{R}$, as follows
@@
$\begin{array}[]{c}\displaystyle a(\bm{\sigma},\bm{\tau}):=\int_{\Omega}\mathcal{C}^{-1}\bm{\sigma}:\bm{\tau},\quad b(\bm{\tau},(\mathbf{v},{\bm{\eta}})):=\int_{\Omega}\mathbf{v}\cdot\mathbf{div}\,\bm{\tau}+\int_{\Omega}\bm{\tau}:{\bm{\eta}},\quad c(\bm{\tau},q):=\dfrac{\alpha}{d\lambda+2\mu}\int_{\Omega}q\,\mathop{\rm tr}\bm{\tau},\\[12.91663pt] \displaystyle\widetilde{a}_{\widehat{\bm{\sigma}},\widehat{p}}({\bm{\varphi}},{\bm{\psi}}):=\int_{\Omega}{\bm{\kappa}}(\widehat{\bm{\sigma}},\widehat{p})^{-1}{\bm{\varphi}}\cdot{\bm{\psi}},\quad\widetilde{b}({\bm{\psi}},q):=\int_{\Omega}q\mathop{\rm div}\nolimits{\bm{\psi}},\quad\widetilde{c}(p,q):=\big{(}c_{0}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}\big{)}\int_{\Omega}p\,q,\end{array}$ (2.1)
@@
respectively, and linear functionals $H\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)^{\prime}$, $F\in(\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega))^{\prime}$, $\widetilde{H}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)^{\prime}$, ${\tilde{F}}\in\mathrm{L}^{2}(\Omega)^{\prime}$
@@
$H(\bm{\tau}):=\langle\bm{\tau}\mbox{\boldmath{$n$}},{\bm{u}}_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}},\quad F(\mathbf{v},{\bm{\eta}}):=\displaystyle-\int_{\Omega}\bm{f}\cdot\mathbf{v},\quad\widetilde{H}({\bm{\psi}}):=\langle{\bm{\psi}}\cdot\mbox{\boldmath{$n$}},p_{\mathrm{D}}\rangle_{\Gamma_{\mathrm{D}}},\quad{\tilde{F}}(q):=\displaystyle-\int_{\Omega}g\,q,$
@@
we arrive at: find $(\bm{\sigma},{\bm{u}},\bm{\rho},{\bm{\varphi}},p)\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)\times\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega)\times\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)\times\mathrm{L}^{2}(\Omega)$, such that:
@@
$\displaystyle a(\bm{\sigma},\bm{\tau})$ $+\quad\displaystyle b(\bm{\tau},({\bm{u}},\bm{\rho}))$ $+\quad c(\bm{\tau},p)$ $=\displaystyle H(\bm{\tau})$ $\forall\,\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega),$ (2.2a)
@@
$\displaystyle b(\bm{\sigma},(\mathbf{v},{\bm{\eta}}))$ $=\displaystyle F(\mathbf{v},{\bm{\eta}})$ $\forall\,\mathbf{v}\in\mathbf{L}^{2}(\Omega),\,\forall\,{\bm{\eta}}\in\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),$ (2.2b)
@@
$\displaystyle\widetilde{a}_{\bm{\sigma},p}({\bm{\varphi}},{\bm{\psi}})$ $+\quad\displaystyle\widetilde{b}({\bm{\psi}},p)$ $=\displaystyle\widetilde{H}({\bm{\psi}})$ $\forall\,{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega),$ (2.2c)
@@
$\displaystyle\widetilde{b}({\bm{\varphi}},q)$ $-\quad\widetilde{c}(p,q)$ $-\quad c(\bm{\sigma},q)$ $=\displaystyle{\tilde{F}}(q)$ $\forall\,q\in\mathrm{L}^{2}(\Omega).$ (2.2d)
In what follows, we stress that bilinear forms without tildes will refer to the perturbed saddle-point problem of the solid, and bilinear forms with tildes relate with the perturbed saddle-point sub-system of the interstitial fluid. Functions with hats will typically denote fixed quantities (which will be of importance in the fixed-point setting).
@@ 2.2 Stability properties and suitable inf-sup conditions
For the sake of the analysis, we allow the permeability ${\bm{\kappa}}(\bm{\sigma},p)$ to be anisotropic but still require ${\bm{\kappa}}(\bm{\sigma},p)^{-1}$ to be uniformly positive definite in $\mathbb{L}^{\infty}(\Omega)$ and Lipschitz continuous with respect to $p\in\mathrm{L}^{2}(\Omega)$. That is, there exist strictly positive constants $\kappa_{1},\kappa_{2}$ such that
$\kappa_{1}|\mathbf{v}|^{2}\leq\mathbf{v}^{\tt t}{\bm{\kappa}}(\cdot,\cdot)^{-1}\mathbf{v},\qquad\|{\bm{\kappa}}(\cdot,p_{1})^{-1}-{\bm{\kappa}}(\cdot,p_{2})^{-1}\|_{\mathbb{L}^{\infty}(\Omega)}\leq\kappa_{2}\|p_{1}-p_{2}\|_{0,\Omega},$ (2.3)
for all $\mathbf{v}\in\mathbb{R}^{d}\setminus\{\bm{0}\}$, and for all $p_{1},p_{2}\in\mathrm{L}^{2}(\Omega)$.
@@
$\big{|}a(\bm{\sigma},\bm{\tau})\big{|}\leq\dfrac{1}{\mu}\|\bm{\sigma}\|_{\mathbf{div};\Omega}\|\bm{\tau}\|_{\mathbf{div};\Omega},\qquad\big{|}b(\bm{\tau},(\mathbf{v},{\bm{\eta}}))\big{|}\leq\|\bm{\tau}\|_{\mathbf{div};\Omega}(\|\mathbf{v}\|_{0,\Omega}+\|{\bm{\eta}}\|_{0,\Omega}),$ (2.4a)
$\big{|}c(\bm{\tau},q)\big{|}\leq\gamma\|\bm{\tau}\|_{\mathbf{div};\Omega}\|q\|_{0,\Omega},$ (2.4b)
$\big{|}\widetilde{b}({\bm{\psi}},q)\big{|}\leq\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega}\|q\|_{0,\Omega},\qquad\big{|}\widetilde{c}(p,q)\big{|}\leq\widetilde{\gamma}\|p\|_{0,\Omega}\|q\|_{0,\Omega},$ (2.4c)
@@
$\gamma:=\dfrac{\alpha\sqrt{d}}{d\lambda+2\mu}{\quad\hbox{and}\quad}\widetilde{\gamma}:=c_{0}+\dfrac{d\alpha^{2}}{d\lambda+2\mu}.$ (2.5)
@@
$\big{|}H(\bm{\tau})\big{|}\leq\|{\bm{u}}_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}\|\bm{\tau}\|_{\mathbf{div};\Omega},\qquad\big{|}F(\mathbf{v},{\bm{\eta}})\big{|}\leq\|\bm{f}\|_{0,\Omega}\|\mathbf{v}\|_{0,\Omega}\leq\|\bm{f}\|_{0,\Omega}(\|\mathbf{v}\|_{0,\Omega}+\|{\bm{\eta}}\|_{0,\Omega}),$
$\big{|}\widetilde{H}({\bm{\psi}})\big{|}\leq\|p_{\mathrm{D}}\|_{1/2,\Gamma_{\mathrm{D}}}\|{\bm{\psi}}\|_{\mathop{\rm div}\nolimits;\Omega},\qquad\big{|}{\tilde{F}}(q)\big{|}\leq\|g\|_{0,\Omega}\|q\|_{0,\Omega}.$
@@
$\mathbb{V}$ =$=\displaystyle\Big{\{}\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega):\quad\mathbf{div}\,\bm{\tau}=\bm{0}{\quad\hbox{and}\quad}\bm{\tau}=\bm{\tau}^{\tt t}{\quad\hbox{in}\quad}\Omega\Big{\}},$ (2.6a)
$\mathbf{V}$ =$=\Big{\{}{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega):\quad\mathop{\rm div}\nolimits\,{\bm{\psi}}=0\quad\text{in }\,\Omega\Big{\}}.$ (2.6b)
@@
$a(\bm{\tau},\bm{\tau})\geq c_{a}\|\bm{\tau}\|_{\mathbf{div};\Omega}^{2}\quad\forall\,\bm{\tau}\in\mathbb{V}.$ (2.7)
Remark 2.1.
From [30, Lemma 2.2] and [31, Lemma 2.2], we have that the ellipticity constant $c_{a}$ has the form $c_{a}=\hat{c}\frac{1}{\mu}$, where $\hat{c}$ depends on $\Gamma_{\mathrm{N}}$, $|\Omega|$, and the Poincaré constant. Based on the above, $c_{a}$ has a non-zero lower bound.
@@
$\sup_{\bm{0}\neq\bm{\tau}\in\mathbb{H}_{\mathrm{N}}(\mathbf{div};\Omega)}\frac{b(\bm{\tau},(\mathbf{v},{\bm{\eta}}))}{\|\bm{\tau}\|_{\mathbf{div};\Omega}}$ $\geq\beta(\|\mathbf{v}\|_{0,\Omega}+\|{\bm{\eta}}\|_{0,\Omega})\quad\forall\,(\mathbf{v},{\bm{\eta}})\in\mathbf{L}^{2}(\Omega)\times\mathbb{L}_{\mathrm{skew}}^{2}(\Omega),$ (2.8a)
$\displaystyle\sup_{\bm{0}\neq{\bm{\psi}}\in\mathbf{H}_{\mathrm{N}}(\mathop{\rm div}\nolimits;\Omega)}\frac{\widetilde{b}({\bm{\psi}},q)}{\|\psi\|_{\mathop{\rm div}\nolimits;\Omega}}$ $\geq\widetilde{\beta}\|q\|_{0,\Omega}\quad\forall\,q\in\mathrm{L}^{2}(\Omega).$ (2.8b)
@@
$\widetilde{c}(q,q)\geq\widetilde{\gamma}\,\|q\|_{0,\Omega}^{2}.$ (2.9)
Remark 2.2.
Due to the careful choice of the bilinear form $a(\cdot,\cdot)$, constants in continuity estimates (2.4a) –(2.4c) and inf-sup conditions (2.8a) and (2.8b) do not blow up when $\lambda\to\infty$ and $c_{0}\to 0$. In particular, the constants $\gamma$ and $\widetilde{\gamma}$, which depend on $\lambda$ and $c_{0}$ (cf. (2.5)), remain bounded.